OpenStudy (anonymous):

Show all work. Mary purchased a package of 18 different plants, but she only needed 12 plants for planting. In how many ways can she select the 12 plants from the package to be planted?

OpenStudy (anonymous):

so, the question is, how many different ways can we select 12 plants out of a total of 18 plants. Correct?

OpenStudy (anonymous):

That means, how many different combinations are there to select 12 plants out of 18 plants. The answer is there are 18 C 12 ways in which you can select 12 plants out of 18 plants.

OpenStudy (anonymous):

not 12!

OpenStudy (anonymous):

\[\dbinom{18}{12}\]

OpenStudy (anonymous):

3C2 ways, which is essentially 3 ways. you can select plant 1 and plant 2 you can select plant 1 and plant 3 you can select plant 2 and plant 3

OpenStudy (anonymous):

if the above statement is there would be 3 possible orders from choosing 2 from a total of 3 plants

OpenStudy (anonymous):

is true* (typo)

OpenStudy (anonymous):

@ rubin, that is correct.

OpenStudy (anonymous):

\[=\dbinom{18}{6}=\frac{18\times 17\times 16\times 15 \times 14 \times 13}{6\times 5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

\[=17\times 2\times 3\times 14\times 13=7514\]

OpenStudy (anonymous):

the answer is certainly not 18

OpenStudy (anonymous):

you have 18 plant out of which you are choosing 12. this is 18 choose 12 written in math as \[\dbinom{18}{12}\]

OpenStudy (anonymous):

yeah. my bad :(

OpenStudy (anonymous):

the number of ways you can choose 12 things from a set of 18 the formula is \[\dbinom{18}{12}=\frac{18!}{12!(18-12)!}\]

OpenStudy (anonymous):

this is \[\frac{18!}{12!6!}\]

OpenStudy (anonymous):

which = 18564

OpenStudy (anonymous):

so the equation to find this is: (total amount to chose from)/((difference between the two numbers)! x (amount chosen)! ?

OpenStudy (anonymous):

yes you are right it is 18564